OpenStudy (anonymous):

. Divide (5a^2−14a−3)÷(a−3) A. a + 5 B. 5a + 1 C. 2a - 4 D. 5a -4

ganeshie8 (ganeshie8):

familiar wid synthetic division ?

OpenStudy (anonymous):

no :/

ganeshie8 (ganeshie8):

okay, another way is to factor the numerator

OpenStudy (anonymous):

\[ (5a^2−14a−3)÷(a−3) =\frac{ (5a^2−14a−3)}{(a−3)}\] \[=\frac{ (5a^2−15a+a−3)}{(a−3)}=\frac{ [5a(a−3)+1(a−3)]}{(a−3)}\] \[=\frac{ (5a+1)(a−3)}{(a−3)} =(5a+1)\] Thus (5a+1) is the simplified form of the given fraction. @LaBlueEyes2013

ganeshie8 (ganeshie8):

can u factor the numerator ?